What is the y intercept of the line through (1,5) and (3,-4) when the slope is
-4 and 1/2?

Answer by Alan3354(69443)

What is the y intercept of the line through (1,5) and (3,-4) when the slope is
-4 and 1/2?
-------
slope m = 9/2
Use y = mx + b and either point to find b, the y-int
---
5 = (9/2)*1 + b
b = 1/2